the sum of the second and third terms of a G.P is six times the fourth term. find the two possible values of the common difference.
Answer by amarjeeth123(569)

Let the first term of the GP be a.
Let the common ratio be r.
second term=ar
third term=ar^2
fourth term=ar^3
We have ar+ar^2=6ar^3
1+r=6r^2
6r^2-r-1=0
6r^2-3r+2r-1=0
3r(2r-1)+1(2r-1)=0
(3r+1)(2r-1)=0
r=-1/3 or r=1/2
The two possible values of the common difference are r=-1/3 and r=1/2.